Skip to content

OSG Technology Area Meeting, 5 April 2022


  • TimT OOO this week
  • BrianL largely unavailable next week due to ITLP workshop
  • Mat OOO Apr 15-29
  • Next week's meeting canceled

Triage Duty

  • This week: BrianL
  • Next week: Mat
  • 16 (-2) open FreshDesk tickets
  • 0 (+0) open GGUS ticket

Jira (as of Monday)

# of tickets Δ State
182 +3 Open
40 -2 Selected for Dev
30 +1 In Progress
13 -1 Dev Complete
9 -4 Ready for Testing
0 +0 Ready for Release

OSG Software Team

  • Kubernetes hackathon today:
    • Dex for Tiger
    • Update CHTC cache/origin
    • Upgrade production GLOW frontend
  • AI (Mat): increase default key size in renew proxy (SOFTWARE-5112)
  • AI (Mat): document xcache (and xrootd?) upgrade to 3.6 (SOFTWARE-5109)
  • AI (BrianL): python3-rrdtool for EL7 (SOFTWARE-5080)
  • AI (BrianL): update VOMS clients default proxy key bit length to 2048 bits (SOFTWARE-4889)


  • Marco: adding small requested changes to 3.7.7; 3.9.5 in progress, including the code to generate scitokens and wlcg tokens using an external script

  • BrianB ran into an issue with an expired GLOW proxy causing a failure at the factory, even though token auth was enabled; he will send Marco the configuration that caused it

  • CMS is still sending 1024-bit proxies which no longer work on EL 8. We need to find out where their proxies are generated and ask them to upgrade. EL 8 can be configured to accept 1024-bit proxies again, but this makes sites less secure so we do not want to recommend it

  • New security release of Go upcoming; no details on the issue yet, but we may need to rebuild our software if it's affected

  • CMS has requested that we build worker node images using Alma Linux, because they have not validated CentOS Stream 8. (WLCG sites in general have been avoiding Stream 8 because it goes EOL in the middle of Run 3.) OS SW does not currently test Alma but that would not be too difficult to add

Support Update

  • OSG Connect (BrianL): help with AP update to OSG 3.6
  • UNL (Derek): Correct whole node GRACC accounting

  • UNL (Derek): help with GRACC records issues. Since UNL is running OSG 3.5 for now, they can use the condor batch system probe

  • CRIC (Mat): with BrianB's help found a Topology issue that was causing SSL errors for CRIC, caused by CentOS Stream 8 dropping a needed package from their default container image

OSG DevOps

  • Go update is coming thursday, will be watching for possible impact on OSG things. Security team is notified.
  • Working on shoveler docs


None this week

OSG Release Team

No updates this week


None this week